OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 610! If four times the number of one-step equations minus twice the number of two-step equations is equal to 340, how many one-step equations auditioned today? 260 340 350 670

OpenStudy (anonymous):

This answer will only tell you the number of one-step equations that have been eliminated. We don't actually know how many auditioned. Also, be careful with the use of ! in mathematics as 1070! = 1070 x 1069 x 1068 x.......x 3 x 2 x1 Let A be the number of one-step equations eliminated and B the number of two-step equations eliminated then, A + B = 1070 .....Equation (1) and we are told that, 3A - 2B = 110.....Equation (2) multiplying Eqn (1) by 2 gives 2A + 2B = 2140........now add Eqn(2) 3A - 2B = 110 5A = 2250 ⇒ A = 450....which is the number of one-step equations ELIMINATED today. http://answers.yahoo.com/question/index?qid=20110903113658AANB9Kr

OpenStudy (anonymous):

260,believe me

OpenStudy (anonymous):

how plss elaborate @lbf

OpenStudy (anonymous):

can i get a medal for that effort

OpenStudy (anonymous):

@Rohangrr 450 is not part of my options

OpenStudy (anonymous):

his answer:A + B = 1070 .....Equation (1) and we are told that, 3A - 2B = 110.....Equation (2) mine: A+B=610;4A-2B=340;DID HE REALLY READ THE PROBLEM?